# Img2CAD: Reverse Engineering 3D CAD Models from Images through VLM-Assisted Conditional Factorization

### v1

@@ 4.2. TrAssembler – Continuous Attribute Prediction in Shared Space
Our exploration led to the finding that, despite GPT-4V showing promise in inferring the discrete base structure – part semantics and their underlying sequence of command types – it exhibits limitations in accurately predicting the continuous attributes of CAD commands. Figure 6 illustrates a failure case of GPT-4V. To address this shortcoming, we introduce TrAssembler, which takes in a single image and, conditioned on the discrete base structure from GPT-4V, predicts the continuous attributes for all CAD commands.
Specifically, TrAssembler is a transformer-based architecture that conditions on the base structure from GPT-4V, represented as a sequence of commands $[C^{j}_{k}]$ for $k\in 1,2,\ldots,N_{j}$ and a semantic part comment $y_{j}$ for each part $P_{j}$ for all parts, i.e., $j\in 1,2,\ldots,M$, where $M$ is the number of parts in GPT-4V’s predicted structure and $N_{j}$ is the number of CAD commands for part $P_{j}$. TrAssembler then predicts the continuous attribute values for each command $C^{j}_{k}$ for all parts $P_{j}$. Since the underlying continuous attribute values differ for each CAD command type (line, arc, circle, extrude), similar to (Wu et al., 2021), we unify them into a common output space that can be predicted by a shared network structure.
@@
Figure 6. Illustration of GPT-4V’s limitations in predicting continuous attributes. While GPT-4V effectively decomposes and labels parts, it struggles with accurate prediction of continuous attributes such as the leg’s scale and backrest’s extrusion direction (marked in red).
Architecture Details.
Figure 5 shows the network architecture of TrAssembler. Each part $P_{j}$, including the CAD structure or command types along with the part comments, is first fed into a part transformer encoder. This encoder transforms each command type into latent embeddings using a learnable dictionary and performs multi-head self-attention on the sequence of commands to obtain a corresponding sequence of command tokens $\{W_{k}^{j}\mid W_{k}^{j}\in\mathbb{R}^{256}\}$. These command tokens are aggregated into a single part token $T_{j}\in\mathbb{R}^{256}$ via summation, which serves as the learned vector for each discrete part from GPT-4V. Since assembling a shape requires understanding the correlation and interaction between different parts, TrAssembler employs a global transformer encoder that conducts global attention across all parts. Specifically, each learned part token $T_{j}$ is concatenated with an image latent embedding and part comment $y_{j}$ embedding for global structure information through the image, and for semantic information through the part comment, respectively. Both the image and the semantic part comment are encoded with CLIP (Radford et al., 2021). The global transformer encoder outputs a list of refined part tokens $\{\hat{T}_{j}\mid\hat{T}_{j}\in\mathbb{R}^{256}\}$ that are semantically-informed and global structure-aware. Finally, our part transformer decoder concatenates the refined part token $\hat{T}_{j}$ to each command token $W_{k}^{j}$ for the sequence of commands for each part $P_{j}$. It then performs multi-head self-attention across all commands in $P_{j}$, which is subsequently fed into an MLP to predict the continuous attribute values for each command.
@@
A core innovation of TrAssembler lies in its utilization of the discrete structure provided by GPT-4V. By conditioning on this discrete structure, the continuous attribute prediction becomes significantly easier, as the image-to-CAD task conditioned on the discrete base structure simplifies into a continuous problem. Also critical to our TrAssembler is its hierarchical transformer design, comprising the part and global transformers, and its use of consistent semantic labels from GPT-4V. These elements facilitate information sharing across the attribute space, which our experiments show is essential for achieving better performance. Additionally, our design choice of using transformer-based architectures enables the handling of varying numbers of parts $M$ of shapes and different numbers of commands $N_{j}$ for parts. This flexibility is crucial for accommodating the diverse and complex nature of CAD tasks.
@@
Ours 0.1685 81.79 73.32 0.1655 83.75 69.55 0.2327 66.32 57.33
Figure 7. Data samples from our curated CAD dataset. For each pair, left: ShapeNet model, right: our CAD-ified model.
@@ 5. Dataset
We need to train TrAssembler with the ground-truth attribute parameters for each part. Therefore, a substantial, curated CAD dataset is necessary. Previous resources such as the ABC dataset (Koch et al., 2019) offer extensive CAD data; however, these datasets predominantly focus on industrial components and are less applicable to domestic items like chairs and tables. To address this gap and further facilitate CAD research on everyday objects, we have compiled a CAD-ified dataset of common objects encompassing three classes from ShapeNet (Chang et al., 2015): chair, table, and cabinet.
For dataset creation, we employ a hybrid method that capitalizes on GPT-4V’s structural priors in a semi-automatic fashion. Initially, we input segmented per-part images (since we have the ground-truth segmentation) into GPT-4V, which then predicts the semantic label and the corresponding lines, arcs, and circles for each part. Subsequently, utilizing the ground-truth labels, we perform one-to-one part label matching using the linear sum assignment method (Burkard et al., 2012). For each matched part prediction, we adjust the ground-truth 3D mesh to correct the scale and translation errors introduced by GPT-4V’s assembly inaccuracies. Adjustments are made to sketch endpoints and extrusion distances in order to align with the ground-truth scale, while the extrusion plane’s origin is repositioned for accurate translation. Additionally, arcs and circles require special treatment for scaling, as direct scaling could deform a circle into an oval, which is not feasible in standard CAD commands. To overcome this, we employ the RANSAC method (Fischler and Bolles, 1981) to accurately fit the parameters for arcs and circles. This hybrid approach has proven to yield superior data quality.
Post data generation, each example undergoes manual review to ensure high fidelity, culminating in a comprehensive dataset comprising 1,026 chairs, 3,243 tables, and 305 cabinets. We follow PartNet (Mo et al., 2019b) train/val/test splits. To the best of our knowledge, this is the first large-scale dataset for common domestic objects, and it is intended to serve as a benchmark for training and evaluating CAD-related tasks on common objects.
@@ 6. Experiments
Ours 0.1685 81.79 73.32 0.1655 83.75 69.55 0.2327 66.32 57.33
w/ Quantized Loss 0.1912 79.82 73.22 0.1881 81.77 67.40 0.3169 63.89 56.12
w/o Hierarchy 0.3282 64.64 49.90 0.2670 76.53 56.07 0.3013 62.99 54.77
w/o Semantics 0.2847 63.14 44.06 0.2645 49.10 32.17 0.2947 59.78 50.21
Baselines.
@@
  • GPT-Only: This baseline directly employs the output CAD program from GPT-4V without our attribute prediction network.
  • Image-3D-CAD: Since most CAD-related works require a 3D mesh as input, we build this baseline to test how well current CAD reconstruction methods perform when converting a 2D image to 3D. We first use Point-E (Nichol et al., 2022) to predict the corresponding point cloud of the input image. Then, we employ a neural Galerkin solver (Huang et al., 2022) to reconstruct the surface, forming a watertight mesh. Finally, the mesh is sent to a self-supervised SECAD-Net (Li et al., 2023) to predict the final CAD program.
  • DeepCAD: We load the checkpoint of DeepCAD pretrained on the ABC dataset (Koch et al., 2019) and finetune it on our dataset. To enable image-to-CAD reconstruction, we follow the original approach in Wu et al. (2021), which involves a second-stage training to align the image encoder with the program encoder by minimizing the L2 loss over the latent vectors while keeping the decoder frozen.
  • DeepCAD-End2End: We also include a variant of DeepCAD that jointly finetunes the image encoder and program decoder in an end-to-end manner. Empirically, we find this approach generally outperforms the naive DeepCAD, as it allows for the joint optimization of the image encoder and program decoder directly under the image-to-CAD prediction objective.
Evaluation metrics.
@@
  • Part Segmentation Accuracy (Seg Acc %): This measures how semantically close the reconstructed object is to the ground truth, specified as the part segmentation accuracy over the point cloud sampled on the surface.
@@
  • Part Segmentation mIoU (Seg mIoU %): This metric also assesses semantic closeness, specified as the mean Intersection over Union (mIoU) of the part segmentation over the point cloud sampled on the surface. We leverage PointNet (Qi et al., 2017) pretrained on ShapeNet (Chang et al., 2015) to produce the segmentation from the generated CAD models. We then compare these segmentations with the ground-truth segmentation labels from ShapeNet to obtain Seg Acc and Seg mIoU. More details can be found in our supplementary material.
Main results.
@@
We present the image-to-CAD reconstruction results of our method against the baselines in Table 1. Qualitative results are illustrated in Figure 9. Img2CAD consistently outperforms the baselines across all three object categories and metrics, demonstrating its strong capability in producing both geometrically and structurally accurate CAD models from image inputs. Notably, Img2CAD reduces the Chamfer distance from 0.3698 to 0.1655 and improves segmentation accuracy and mIoU by 16.38% and 17.34%, respectively, over the best-performing baseline, DeepCAD-End2End, on the Table category.
@@
The GPT-Only baseline performs significantly worse than the other approaches, validating that GPT-4V alone is insufficient to predict both the program structure and continuous attributes directly. This supports our approach of conditionally factorizing the task: using the program structure predicted by GPT-4V and offloading the attribute prediction to a learned network. Both DeepCAD and DeepCAD-End2End achieve better performance than GPT-Only but are still markedly inferior to Img2CAD.
Table 3. Ablation study of GPT prompts on Chair test dataset.
Reminder ContExample CoT Prog Succ $\uparrow$ CD $\downarrow$ Part IoU $\uparrow$
- - - 89.02 0.1601 59.82
$\checkmark$ - - 96.34 0.1388 68.22
$\checkmark$ $\checkmark$ - 98.78 0.1349 68.59
$\checkmark$ $\checkmark$ $\checkmark$ 98.78 0.1248 73.59
Ablation on TrAssembler.
@@
We ablate several core designs of our TrAssembler with results depicted in Table 2 and qualitatively visualized in Figure 10. Note that these variants are all conditioned on the program structure predicted by GPT-4V.
We first implement a variant of TrAssembler that quantizes the continuous parameters into 256 bins, turning parameter fitting into a classification problem. This variant generally leads to slightly worse performance compared to TrAssembler, highlighting the advantage of directly predicting continuous values for CAD parameters. Notably, the variant of TrAssembler with quantized loss still outperforms DeepCAD, showcasing the benefit of conditional factorization – i.e., using GPT-4V for discrete structure and a learned network for continuous parameters.
We further investigate the importance of using both (i) hierarchical transformer design and (ii) semantic part comment information in our TrAssembler. w/o Hierarchy removes the hierarchical design but retains the semantic labels. Specifically, it concatenates part label embeddings and command tokens into a 1D array that is fed through a single, flat transformer. This variant demonstrates how essential the hierarchical structure is for effectively capturing both local and global interactions within the CAD model, as the performance drops significantly without it. w/o Semantics keeps the hierarchical design but excludes the semantic part comments, showing that even with a robust hierarchical architecture, the absence of semantic information results in poorer performance.
These results, depicted in Table 2 and visualized in Figure 10, underscore the necessity of both the hierarchical transformer design and the inclusion of semantic part comments for optimal performance in TrAssembler.
Ablation studies on GPT-4V prompt.
We further discuss the importance of three techniques we adopt in GPT-4V prompting:
  • Reminder: We add several bullet-point reminders to ensure the CAD program is complete without missing any parts (e.g., ”Give the FULL PROGRAM for ALL PARTS”).
  • In-context Example (ContExample): We provide GPT-4V with an example program for a chair’s backrest.
  • Chain of Thoughts (CoT): We instruct GPT-4V to first enumerate all object parts in the image and then output programs for each part, rather than generating a single full program.
For this ablation, we use the oracle assembling process (discussed in Section 5) instead of TrAssembler to eliminate any confounding factors introduced by the assembling process. We report CD together with two additional metrics:
  • Program Success rate (Prog Succ %): This reflects how effectively GPT-4V generates a valid CAD program.
  • Part IoU (%): This measures the consistency of GPT-generated semantic part labels with the ground-truth labels. Part IoU is computed by measuring the average frequency with which the predicted label and its corresponding ground-truth label have a CLIP embedding distance larger than 0.8.
The results of these ablation studies, detailed in Table 3, demonstrate the critical role of these prompting techniques in improving the quality and completeness of the generated CAD programs. The inclusion of reminders, in-context examples, and chain of thought reasoning significantly enhances both the program success rate and the part IoU, underscoring the necessity of these strategies for effective GPT-4V prompting. More details on these prompt modifications and metrics can be found in the supplementary materials.
Robustness on Arbitrary Viewpoints.
@@
Frontal View 0.1889 77.29 66.73
Arbitrary View 0.1941 76.65 65.95
Our method is also very robust thanks to the view invariance provided by large vision foundation models (i.e., GPT-4V and CLIP). As shown in Table 4, we achieve almost the same performance when evaluated on images sampled from arbitrary viewpoints. This important property helps make our method generalizable to images in the wild, even when it is trained on frontal-view images. Qualitative examples can be found in Figure 8.
Figure 8. Example output visualizations given arbitrary-viewpoint input images.

### v2

@@ 4.2. TrAssembler – Continuous Attribute Prediction in Shared Space
Our exploration shows that despite VLM being good at inferring the discrete base structure, part semantics and their underlying sequence of command types, it exhibits limitations in accurately predicting the continuous attributes of CAD commands. Figure 6 illustrates a failure case of VLMs, even with Llama3.2 finetuned on both structure and continuous attributes. To address this, we introduce TrAssembler, which takes in a single image and conditioned on the discrete base structure from our finetuned VLM (4.1), uses flow matching to denoise noisy continuous attributes $\mathbf{x}_{t}$ for all CAD commands. $t$ is the denoising timestep.
TrAssembler first encodes each part by processing its command sequence and noisy attribute inputs through a dedicated part transformer, incorporating cross-attention with DINOv2 image features to produce part-level tokens. These tokens are concatenated with CLIP-based semantic embeddings and globally refined via a transformer to capture inter-part relationships. Finally, a decoder predicts denoising flows for each command attribute by cross-attending to learnable parameter tokens and regressing flow scalars via an MLP. The overall architecture of TrAssembler is illustrated in Figure 5.
To sample continuous CAD attributes, we employ GMFlow (Chen et al., 2025b) as our flow matching generative model (Lipman et al., 2023; Liu et al., 2022), which predicts a dynamic Gaussian mixture distribution $q(\mathbf{u}|\mathbf{x}_{t})$ of flow velocity $\mathbf{u}$ conditioned on noisy attributes $\mathbf{x}_{t}$ at diffusion time $t$. The model is trained with the negative log-likelihood loss on flow velocity:
(1) $\mathcal{L}=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\left[-\log q\left(\frac{\mathbf{x}_{t}-\mathbf{x}_{0}}{t}\middle|\mathbf{x}_{t}\right)\right],$
where $\mathbf{x}_{0}$, $\mathbf{x}_{t}$ are ground truth and noisy attributes, respectively. Compared to standard flow matching, GMFlow better captures multi-modal characteristics of CAD attributes. Please see supplementary for more details on our net architecture and GMFlow.
@@
Figure 6. Illustration of VLM’s limitations in predicting continuous attributes. While VLMs effectively decomposes labeled parts, it struggles with predicting accurate continuous attributes.:
@@
A core innovation of TrAssembler lies in its utilization of the discrete structure provided by VLM. By conditioning on this discrete structure, the continuous attribute prediction becomes significantly easier, as the image-to-CAD task conditioned on the discrete base structure simplifies into a continuous problem. Also critical to our TrAssembler is its hierarchical transformer design, comprising the part and global transformers, and its use of consistent semantic labels from VLM. These elements facilitate information sharing across the attribute space, which our experiments show is essential for achieving better performance. Additionally, our design choice of using transformer-based architectures enables the handling of varying numbers of parts and different numbers of commands.
4.3. Inference Time Symmetry Guidance
Many human-made objects exhibit strong symmetry, a property often overlooked by standard flow-matching or diffusion methods. To address this, we introduce symmetry score guidance during each inference ODE solver step of GMFlow. After each inference step, we refine $\mathbf{x}_{t}$ by performing gradient descent through our estimated score function over the symmetric data manifold:
(2) $\mathbf{x}_{t}=\mathbf{x}_{t}-\lambda\nabla\log p_{\text{sym}}(t)=\mathbf{x}_{t}-\lambda\frac{\partial\mathcal{L}_{\text{sym}}}{\partial\mathbf{x}_{t}},$
where $\mathcal{L}_{\text{sym}}$ denotes the symmetry loss. This loss is computed by sampling a set of points, reflecting them across a symmetry plane or rotating them according to a detected axis of rotational symmetry, and then measuring the absolute differences between the signed distance functions (SDFs) of the original and transformed sets. This guidance enforces symmetry manifold constraint on the CAD attributes $\mathbf{x}_{t}$. Additionally, we prompt and query GPT-4o to identify the specific type of symmetry present in the input image, e.g., planar or rotational symmetry. In this work, we optimize for a single symmetry type per shape, which we find to be sufficient for producing consistent improvements.
@@ 4.3. Inference Time Symmetry Guidance
Ours 0.0984 91.86 87.88 0.0966 93.81 82.13 0.1573 73.08 57.59
Figure 7. Data samples from our curated CAD dataset. For each pair, left: ShapeNet model, right: our CAD-ified model.:
@@ 5. Dataset
We need to train TrAssembler with the ground-truth attribute parameters for each part. Therefore, a substantial, curated CAD dataset is necessary. Previous resources such as the ABC dataset (Koch et al., 2019) offer extensive CAD data; however, these datasets predominantly focus on industrial components and are less applicable to domestic items like chairs and tables. To address this gap and further facilitate CAD research on everyday objects, we have compiled a CAD-ified dataset of common objects encompassing three classes from ShapeNet (Chang et al., 2015): chair, table, and cabinet. We select these categories because their corresponding part semantic labels are annotated in PartNet (Mo et al., 2019b), which are subsequently used in the dataset creation process.
For dataset creation, we employ a hybrid method that capitalizes on GPT-4V’s structural priors in a semi-automatic fashion. First, since we have the ground-truth segmentation, we input the rendered images with segmented parts that we highlight in red (see supplementary for an example) into GPT-4V, which then predicts the semantic label and the corresponding lines, arcs, and circles for each part. Subsequently, utilizing the ground-truth PartNet labels, we perform one-to-one part label matching using the linear sum assignment method (Burkard et al., 2012). For each matched part prediction, we adjust the predicted 3D mesh to correct the scale and translation errors introduced by GPT-4V’s assembly inaccuracies. The resulting adjusted 3D meshes serve as the ground truth labels in our curated dataset. More details can be found in our supplementary.
After data generation, each sample undergoes manual human review to ensure its quality about which 30% are kept), culminating in a comprehensive dataset comprising 1,026 chairs, 3,243 tables, and 305 cabinets. We follow PartNet (Mo et al., 2019b) train/val/test splits. To the best of our knowledge, this is the first large-scale dataset for common domestic objects, and it is intended to serve as a benchmark for training and evaluating CAD-related tasks on common objects.
@@ 6. Experiments
(i) Llama3.2 end-to-end 0.2568 62.98 50.12 0.2432 72.87 61.49 0.3707 59.18 55.11
(ii) Our finetuned VLM + Flat TrAssembler 0.3282 64.64 49.90 0.2670 76.53 56.07 0.3013 62.99 54.77
(iii) Our finetuned VLM + Hierarchical Design 0.2847 63.14 44.06 0.2645 49.10 32.17 0.2947 59.78 50.21
(iv) iii + Semantic Part Info 0.1685 81.79 73.32 0.1655 83.75 69.55 0.2327 66.32 57.33
(v) iv + Flow Matching 0.0992 91.45 86.82 0.0993 94.33 82.49 0.1554 71.81 59.03
(vi) v + Symmetry Guidance (Ours) 0.0984 91.86 87.88 0.0966 93.81 82.13 0.1573 73.08 57.59
6.1. Baselines.
@@
  • GPT-4o: This baseline directly employs GPT-4o to give the CAD program of the input.
  • Image-3D-CAD: Since most CAD-related works require a 3D mesh as input, we build this baseline to test how well current CAD reconstruction methods perform when converting a 2D image to 3D. We first use LRM (Hong et al., 2023) to predict the corresponding mesh of the input image.
    Then, the mesh is sent to a self-supervised SECAD-Net (Li et al., 2023) to predict the final CAD program.
  • DeepCAD: We load the checkpoint of DeepCAD pretrained on the ABC dataset (Koch et al., 2019) and finetune it on our dataset. To enable image-to-CAD reconstruction, we follow the original approach in (Wu et al., 2021), which involves a second-stage training to align the image encoder with the program encoder by minimizing the L2 loss over the latent vectors while keeping the decoder frozen.
  • DeepCAD-End2End: We also include a variant of DeepCAD that jointly finetunes the image encoder and program decoder in an end-to-end manner.
6.2. Evaluation metrics.
@@
  • Part Segmentation Accuracy (Seg Acc %): This measures how semantically close the reconstructed object is to the ground truth, specified as the part segmentation accuracy over the point cloud sampled on the surface. We leverage PointNet (Qi et al., 2017) pretrained on ShapeNet (Chang et al., 2015) to produce the segmentation from the generated CAD models. We then compare these segmentations with the ground-truth segmentation labels from ShapeNet to obtain Seg Acc and Seg mIoU. More details in supplementary.
  • Part Segmentation mIoU (Seg mIoU %): This metric also assesses semantic closeness, specified as the mean Intersection over Union (mIoU) of the part segmentation over the point cloud sampled on the surface.
6.3. Main results.
@@
We present the image-to-CAD reconstruction results of our method against the baselines in Table 1. Qualitative results are illustrated in Figure 8. Img2CAD consistently outperforms the baselines across all three object categories and metrics, demonstrating its strong capability in producing both geometrically and structurally accurate CAD models from image inputs. Notably, Img2CAD reduces the average Chamfer distance from 0.3108 to 0.1174 and improves segmentation accuracy and mIoU by 17.94% and 19.03%, respectively, over the best-performing baseline, DeepCAD-End2End.
@@
GPT-4o performs worse than the other approaches, validating that even state-of-the-art VLMs are insufficient to predict both the program structure and continuous attributes directly. This supports our approach of conditionally factorizing the task: using the program structure predicted by VLMs and offloading the attribute prediction to a learned network. Both DeepCAD and DeepCAD-End2End achieve better performance than GPT-4o but are still markedly inferior to Img2CAD.
Figure 8. Qualitative results between different methods on the test split of ShapeNet CADified dataset. Our method achieves much better results.:
6.4. Ablation studies.
@@
We ablate several core designs of our TrAssembler with results depicted in Table 2 and qualitatively visualized in Figure 9. Note that these variants are all conditioned on the program structure predicted by Llama.
  • (i)
    Llama3.2 end-to-end: First we ablate on the necessity of decomposing the task into predicting the discrete global structure and continuous attributes. Specifically, we compare against a variant in which LLaMA3.2 is directly finetuned to predict both CAD structures and attributes.
  • (ii)
    Flat TrAssembler: We apply conditional factorization by using our finetuned VLM on discrete global structure, but use a flat transformer that treats all attributes from different parts as a single 1D token sequence. These are processed through a non-hierarchical transformer.
  • (iii)
    Hierarchical Design: We apply conditional factorization by using our finetuned VLM on discrete global structure, and further introduce our hierarchical transformer architecture, consisting of the part transformer and the global transformer, without part label information.
  • (iv)
    Semantic Part Info: We then further incorporate semantic information by adding part-specific text embeddings, enabling the model to better reason about part semantics.
  • (v)
    Flow Matching: We replace the naive feed-forward prediction head with our flow matching-based loss.
  • (vi)
    Symmetry Score Guidance: At test time, we further introduce a symmetry-aware guidance strategy to encourage more regular and symmetric outputs.
Quantitative results are reported in Table 2 and qualitative visualizations are shown in Figure 9. Experimental results indicate that this approach fails to generalize well to test-time shapes, primarily due to LLaMA3.2’s limited capacity for reasoning over 3D geometry. Moreover, the results show that the hierarchical transformer design and semantic part comments together improve performance by capturing both local and global part-level interactions. The flow matching strategy provides further gains through a diffusion-inspired structured prediction. Finally, symmetry guidance at inference time helps refine the output structure, better visualized in Figure 9.
Figure 9. Qualitative comparisons of our ablation studies.:
6.5. Symmetry and Connectivity.
To evaluate the structural fidelity of the generated CAD programs, we report two quantitative metrics averaged across all object categories: the number of strongly connected components (SCCs) and the symmetry Chamfer distance. The SCC metric captures geometric disconnections by counting the number of connected components in the output mesh, where two vertices are considered connected if their Euclidean distance is below a threshold of 0.05. Fewer SCCs indicate more coherent and physically plausible structures (fewer floating components). The symmetry Chamfer distance measures how closely the generated shape aligns with its mirror reflection over the X-axis, using bidirectional Chamfer distance between point clouds sampled from the original and mirrored meshes. As shown in Table 3, introducing hierarchical design with semantic part information already yields notable improvements in both connectivity and symmetry. Our full model, equipped with symmetry guidance, further reduces structural fragmentation and enhances global regularity, producing CAD programs that are both functionally and geometrically well-formed.
Table 3. Structural quality analysis: number of strongly connected components (SCC, lower is better) and symmetry chamfer (lower is better).
#SCC $\downarrow$ SymChamfer $\downarrow$
(i) Llama3.2 end-to-end 2.06 0.1268
(ii) Our finetuned VLM + Flat TrAssembler 1.79 0.1502
(iii) Our finetuned VLM + Hierarchical Design 1.90 0.2748
(iv) iii + Semantic Part Info 1.49 0.1176
(v) iv + Flow Matching 1.16 0.0827
(vi) v + Symmetry Guidance (Ours) 1.11 0.0756
6.6. Robustness on arbitrary viewpoints.
@@
Frontal View 0.1174 86.25 75.87
Arbitrary View 0.1396 84.41 73.53
Our method is also very robust thanks to the view invariance provided by VLMs(i.e., Llama, DINO and CLIP). As shown in Table 4, we get a slight drop on performance when evaluated on images sampled from arbitrary viewpoints. This important property helps make our method generalizable to images in the wild, even when it is trained on frontal-view images, as shown in Figure 10. We also computed the Chamfer Distance (CD) between the instance-wise CAD reconstructions from frontal and arbitrary views, obtaining a low CD of 0.0282. Qualitative comparisons between the two views are presented in Figure 11.
Figure 10. Example output visualizations given arbitrary-view inputs.:
Figure 11. Example visualizations of reconstruction results from frontal and arbitrary-view inputs.: